### v2

# Nudging Nutrition: Lessons from the Danish “Fat Tax”

Abstract
In October 2011, Denmark introduced the world’s first and, to date, only tax targeting saturated fat. However, this tax was subsequently abolished in January 2013. Leveraging exogenous variation from untaxed Northern-German consumers, we employ a difference-in-differences approach to estimate the causal effects of both the implementation and repeal of the tax on consumption and expenditure behavior across eight product categories targeted by the tax. Our findings reveal significant heterogeneity in the tax’s impact across these products. During the taxed period, there was a notable decline in consumption of bacon, liver sausage, salami, and cheese, particularly among low-income households. In contrast, expenditure on butter, cream, margarine, and sour cream increased as prices rose. Interestingly, we do not observe any difference in expenditure increases between high and low-income households, suggesting that the latter were disproportionately affected by the tax. After the repeal of the tax, we do not observe any significant decline in consumption. On the contrary, there was an overall increase in consumption for certain products, prompting concerns about unintended consequences resulting from the brief implementation of the tax.
Keywords: Fat tax, consumer responses, difference-in-differences, health policy.
JEL classification: D12, I18, H31, C21.
We express our gratitude for the financial support provided by the Independent Research Fund Denmark, grant 6109-00075B, and Helsefonden, grant 18-B-0130. Additionally, we acknowledge the valuable feedback received from Phillip Heiler, Toru Kitagawa, Michael Lechner, Volha Lazuka, Pedro Sant’Anna, Jeff Smith, and the participants of the 2021 DGPE workshop, as well as the participants of the “Applied Econometrics of Program Evaluation” course. Special thanks are extended to Joachim Dahl Steensbjerre for his exceptional research assistance.
1 Introduction
Cardiovascular diseases are the world’s primary cause of mortality, and their prevalence due to modifiable risk factors like unhealthy diets is rising (Roth et al., , 2020). Since reducing saturated fat intake has been linked to a lower risk of cardiovascular diseases (Hooper et al., , 2020), it is no surprise that the World Health Organization strongly recommended to keep saturated fat content below 10% of the total energy intake (World Health Organization, , 2023). In Denmark, the dietary intake of saturated fats equaled 15% in 2010, surpassing the recommended levels and exhibiting relatively higher values compared to its European counterparts (Eilander et al., , 2015). Specifically, meat was found to account for 19% of the total saturated fat intake, while dairy products contributed with 30%, and fat-based products such as butter and margarine contributed with 32%.
The excessive intake of saturated fat was the main reason for introducing a tax on saturated fat, the so-called “fat tax”, which was in place in Denmark between October 2011 and January 2013. It was a national initiative, applying to foods with saturated fat content exceeding 2.3%, including fats like margarine and butter, meats, and dairy products, but excluding drinking milk and milk-based yogurts. Despite its significance, there is limited empirical evidence on the effectiveness of the Danish fat tax to shape consumer responses, leaving policymakers with little guidance.
The economic rationale behind the implementation of food taxes, which aim at improving the healthiness of diets, is to address the externalities associated with the consumption of unhealthy foods. The effectiveness of a tax hence depends on consumers actually changing behavior in accordance with the tax (Cawley et al., , 2019). Assessing the efficacy of food taxes remains a complex task, influenced by a range of factors that shape consumer behavior and producer responses. Moreover, the overall impact on public health is contingent on consumer decisions regarding available substitutes and the healthiness of these alternatives. A few studies have estimated the effect of fat taxes using food price elasticities to simulate the effect of a fat tax on consumer behavior (Chouinard et al., , 2007; Allais et al., , 2010; Nordström and Thunström, , 2009), but their findings lack a causal interpretation (Powell et al., , 2013). Furthermore, these analyses assume that the tax is fully passed on to the consumers, which might not match reality. Theoretical studies on the consumer responses to food taxes (Lusk and Schroeter, , 2012; Yaniv et al., , 2009; Gortmaker et al., , 2011) stress the need for more robust empirical evidence using real-world data. A limited number of studies evaluated the Danish fat tax using real-world data (Jensen and Smed, , 2013; Jensen et al., , 2016; Bødker et al., 2015a, ; Smed et al., , 2016), resulting in inconclusive findings. In a recent review, Lhachimi et al., (2020) point out that two of these studies (Jensen and Smed, , 2013; Jensen et al., , 2016) lack an analysis of the effects of the tax on household expenditure. More importantly, all of these studies lack a control group that is not exposed to the tax and therefore lack a causal interpretation of the effect of the tax on consumer behavior. This forms a notable gap that is addressed in this study.
Our study design is unique in that we have access to a Northern-German households panel, which we use to construct a control group from a geographic area that is culturally close to Denmark in terms of food consumption behaviors, but is not affected by the tax. This allows us to estimate the causal effect of the tax on consumption and expenditure behaviors for different food groups employing the difference-in-differences (DID) estimator proposed by Callaway and Sant’Anna, (2021). We use Callaway and Sant’Anna, (2021) as our estimator for three reasons: it allows us to control for observed household characteristics in a flexible way; it is doubly robust, and it provides conservative confidence intervals.
We contribute to the existing literature in several ways. First, having access to a well-defined control group adds a causal perspective to the limited literature evaluating the impact of the fat tax on consumption of foods high in saturated fat. Second, we can estimate the causal effect of the tax on expenditure behaviors and the prices actually paid by the consumer. Third, our unique data allows us to study how the effect of the tax evolves over time. In particular, we offer new insights into the aftermath of the abolishment of the tax in January 2013, shedding light on potential habit formation and unintended effects on prices in the post-tax period. Moreover, we are able to estimate the causal effect of the tax on prices for each product group and measure how the tax was actually passed over to consumers. Finally, we estimate effect heterogeneity with respect to income levels to see whether low-income households were more affected by the tax than higher-income households. This is crucial to gain insight into the regressivity of the tax.
Our identification strategy relies on the assumption that purchases in Denmark and Germany would have followed the same trends had the tax not been introduced. Even though this assumption is essentially untestable, we study its plausibility through various pre-trend tests. A second crucial assumption is that there are no anticipation effects. The tax was announced in March 2011 before it was implemented in October 2011. We do not find any empirical evidence of hoarding behaviors in the two quarters prior to the tax implementation. However, to completely exclude the presence of potential anticipation effects, such as hoarding, observations six months before the introduction of the tax are excluded. Furthermore, to strengthen the validity of our results, we perform robustness checks with respect to cross-border shopping by excluding Danish households living close to the border with Germany from the analysis. We further show that our results do not rely on controlling for observed covariates, as they remain virtually unchanged when we use an unconditional DID estimator.
Our main findings suggest that the tax was effective in reducing the consumption of only three of the products we studied: cheese, liver sausage, and salami. This reduction in consumption is only present during the tax period, and we do not find any persistent effects on these products after the tax was abolished. For butter, cream, margarine, and sour cream, we find a significant increase in expenditure, which does not translate into a reduction in consumption. Interestingly, these are the only products for which we find a significant increase in price during the tax period. This price increase persists in the post-tax period for butter and sour cream. For butter, margarine, and sour cream, we observe an overall increase in consumption in the post-tax period. Overall, our findings suggest that the tax has largely failed to induce a long-run reduction in consumption for most products. For butter, cream, margarine and sour cream, the tax actually induced an increase in consumption, at least in the two years following the tax’s abolishment that we are able to observe in our data.
Examining effect heterogeneity between high and low-income households reveals a noteworthy reduction in both consumption and expenditure on cheese and liver sausage, but this reduction is significant only for the latter product group. Furthermore, a significant decrease in bacon consumption is identified exclusively among low-income households. Notably, both high and low-income households exhibited a comparable increase in expenditure on butter and margarine. Finally, the post tax increase, which we find for butter, cream, margarine and sour cream, is largely driven by high-income households.
The remainder of this paper is organized as follows: First, we discuss the institutional background of the Danish fat tax in Section 2. Subsequently, we offer an overview of the data, including details of our identification strategy and its underlying assumptions in Section 3. The results are presented in Section 4, with a focus on each product category individually, along with potential price mechanisms. Section 5 addresses concerns related to our identification strategy, while Section 6 explores potential heterogeneity in tax responses. The paper concludes in Section 7.
2 Background on the Danish fat tax
In October 2011, Denmark introduced a tax targeting saturated fat, with the aim of improving dietary health and with the expectation of raising a tax revenue of 1 billion DKK (156.5 million USD [^3] A conversion rate of 6.39 DKK per USD is used throughout the paper.) (Bødker et al., 2015a, ; Smed, , 2012). The tax amounted to 16 DKK (2.51 USD) per kilo of saturated fat on foods that exceeded a threshold limit of 2.3% saturated fat. This amounts to 20 DKK (3.13 USD) per kilo including the 25% value added tax. The threshold was set such that drinking milk and milk-based yogurts were excluded from the tax. The tax predominantly affected products such as meat, dairy, oil, margarine, butter, and butter blends as these products contain considerable amounts of saturated fat. Producers had the option to use standardized coefficients for meat products instead of the actual fat content in the products. These coefficients were animal-specific (that is, there was one coefficient for pork products, one for beef, etc.), and based on the average content of saturated fat in the various types of meat consumed in Denmark (Smed, , 2012). Generally these coefficients were used instead of specific rates for different meat cuts since the latter option had higher administration costs. The administration of the fat tax was burdensome and estimated to have a retail and whole-sale cost of 200 million DKK (approximately 31 million USD) according to a report by the Danish Chamber of Commerce and the Confederation of Danish Industry (Bødker et al., 2015b, ).
The Danish Tax Committee anticipated an increase in consumer prices up to 7% for meats, up to 33% for dairy, up to 35% for animal fat, up to 16% for cooking oils, up to 24% for margarine and up to 20% for blended butter products (Bødker et al., 2015a, ). Konkurrence- og Forbrugerstyrelsen, (2014) analyzed the price increase of 27 taxed products and found that only for six of these products the consumer prices increased by more than can be explained by increases in the general food price index. They observed an overshifting of the tax, i.e. retailers increased the price on meat by more than what could be justified by the tax, increasing their profit margin by 1% to 5%. This finding is supported by Jensen et al., (2016) who found a price increase of 13% to 16% for high-fat minced beef and cream products. These estimated increases were higher than the a priori expected tax-induced price increases. The same study found small negative effect on prices for lower-fat products (Jensen et al., , 2016).
In an agreement in November 2012, the government decided that the fat tax should be repealed with effect from January 2013. The main reasons stated were increased consumer prices, large administration cost, and the endangerment of Danish jobs (Konkurrence- og Forbrugerstyrelsen, , 2014). [^4] For more information on the Danish tax on saturated fat see Smed, (2012) and Ecorys, Euromonitor and IDEA, DTI, (2014). For more information on the repeal of the Danish tax on saturated fat see Vallgårda et al., (2015).
3 Data, identification strategy, estimation, and inference
3.1 Data
To estimate the impact of the Danish fat tax on consumption behavior, we leverage household panel data from GfK Consumer Scan, which encompasses households located in Denmark as well as households in Northern Germany, specifically Schleswig-Holstein and Hamburg. [^5] For more information on the Danish data set, see Smed, (2008). It is worth noting that these Danish household data are also used in the analysis of Jensen and Smed, (2013) and Smed et al., (2016), yet they lacked the Northern-German households as a control group and the length of the data periods considered differ.
GfK maintains a representative panel with respect to geographic location, education, and age. When a household leaves the panel, it is replaced by a similar household with respect to the aforementioned characteristics. Participants in the panel are equipped with a hand scanning device, enabling them to document detailed information on their weekly grocery shopping. The diary keeper reports the volume of foods purchased for the household, the expenditure, date and time of the purchase, among other details. While surveys and self-reported data are often susceptible to recall bias, this is less likely to be a concern when using scanning devices. A scanning device can ease registration of purchase information, which subsequently might be experienced as less burdensome to the households than registering food consumption manually. Moreover, our data provides an advantage over store scanner data by incorporating purchases made abroad. However, we acknowledge that our data do not account for food acquired outside of stores, such as in restaurants. The purchase diaries are submitted to GfK weekly and thoroughly checked for correct completion and consistency. As highlighted by Smed, (2008), a comparison of the budget-shares of foods from the GfK consumer panel with budget-shares revealed from data of Statistics Denmark shows no major biases for the distribution of expenses between basic food categories, even though the home-scan panel generally under-report in terms of absolute values.
In our analysis, we consider a wide variety of product categories targeted to a different degree by the tax, encompassing bacon, butter and butter blends (referred to as butter), hard cheeses (referred to as cheese), cream, liver sausage and paté (referred to as liver sausage), margarine, salami, and sour cream. The product categories were first and foremost chosen based on their substantial contribution to the saturated fat intake in Denmark. We aggregate different products into products categories in a way that ensures both comparability between Germany and Denmark as well as an adequate number of observations.
Our main analysis uses data covering the period from January 2009 until December 2014. The data are aggregated to a quarterly household panel, encompassing 11 quarters before the tax introduction and 13 quarters after the tax introduction. Furthermore, we create product-specific panels where households that never purchase from the considered product category are excluded. If the number of zero purchases for a given product changes systematically over time differentially between Denmark and Northern Germany, our results could be affected by sample selection bias. However, Figure 9 below shows similar pre-trends in the average number of observations with zero outcome between Denmark and Northern Germany. In general, for all product, we do not find any systematic change in the difference in the number of zeros even in the post-treatment period. The fact that we use quarterly data reduces the number of excluded households and the potential of sample selection bias.
Figure 1: Bacon.
Figure 2: Butter.
Figure 3: Cheese.
Figure 4: Cream.
Figure 5: Liver sausage.
Figure 6: Margarine.
Figure 7: Salami.
Figure 1: Bacon.
Figure 2: Butter.
Figure 3: Cheese.
Figure 4: Cream.
Figure 5: Liver sausage.
Figure 6: Margarine.
Figure 7: Salami.
Figure 8: Sour cream.
Figure 9: The figures show the de-trended quarterly average number of observations with zero value in the outcome variable for Danish and Northern-German households, respectively.
Figures 12 and 13 in Appendix A.1 illustrate the data preparation steps and the number of available observations for each step. Our final analysis incorporates between 2,056 and 2,526 unique Danish households, depending on the product category, and 553 to 1,616 German households.
The data further contain background information on the participating households, which is based on an annual questionnaire filled in by the diary keeper. This includes information on the geographic region of residence, the age of the head of the household, the number of children below age 15, the household size, income level, gender of the diary keeper, and education level. The data set with German households contains similar information as the Danish data. We harmonize the variables that are reported differently in the two data set, such as income and education. [^6] The education level is based on the International Standard Classification of Education (ISCED). To harmonize the income variable, which is reported in different brackets in the two data sets, we create quintiles based on the midpoints of the income ranges in the data set that contains all products for Germany and Denmark. We create five different levels, representing very low income, low income, medium income, high income and very high income. For our analysis, we use the most recently observed pre-tax characteristics for a household. That is, if a household’s income is observed in both the years 2009 and 2010, we use the income level observed in 2010. Pre-treatment characteristics of the households in our study are summarized in Table 6 in Appendix A.2.777Smed et al., (2016) use the same data and compare the descriptive statistics to official numbers from Statistics Denmark. They observe that the panel contains relatively more households in the capital and households with a lower level of educational attainment. There are some notable differences between the Northern-German and Danish households. For instance, the age of the head of the household has a tendency to be higher in Germany compared to Denmark. Additionally, there is a noticeable gender distribution distinction among diary keepers, with 78% being female in Denmark as opposed to only 59% in Germany. Moreover, households in Germany tend to be larger compared to those in Denmark. Since the household characteristics listed in the table are included as control variables in our difference-in-differences strategy, we address the concern that differences in consumption trends between Germany and Denmark might be caused by imbalances in sample composition.
To investigate consumers’ response to the introduction of the fat tax, we define three distinct outcomes of interest. Firstly, we define three variables to investigate consumption and expenditure behavior: (1) the average quarterly household consumption, measured by the average weight or volume in gr or ml purchased per quarter per household, (2) the average number of packages purchased per household, and (3) the average quarterly total expenditure per household in Euro cents. Note that we use product purchases as a proxy for consumption, recognizing that a purchase does not necessarily indicate consumption. To investigate the price changes caused by the tax, we construct a variable for the average quarterly price paid per 100 units (gr or ml depending on the product) in Euro cents, allowing us to investigate whether the tax was passed on to consumers by retailers and producers. [^8] Danish expenditure is converted from DKK to Euro cents using the rate 7.4405. Additionally, since quarters have different number of days we re-weight our outcomes such that each quarter consist of the same amount of days. Finally, we partial out the effect of seasoanl trends from all outcomes. To achieve this we simply run two separate linear regressions, one for each country, for each outcome on quarterly dummies and we use the residuals of these regressions as outcomes. [^9] The pre-tax and tax average quantities of our outcomes of interest are presented in Table 7 in Appendix A.3. Outliers are removed using Tukey’s fences with the constant $k=3$.
3.2 Identification strategy, estimation, and inference
The key strength of our study lies in the ability to observe a proper control group, enabling us to identify a causal effect of the Danish fat tax on consumption behavior. We observe comprehensive information on Northern-German households from Schleswig-Holstein and Hamburg. A suitable control group not only removes time trends, but also unobserved demand shocks (Cawley et al., , 2019). A notable advantage of a geographically close control group is the cultural proximity between the treated and control sample, increasing the likelihood that they are subject to similar external shocks. Our identification strategy relies on the DID framework. The main idea is to compare Danish households to their German counterparts, both before and after the introduction of the fat tax, in terms of our outcome variable and for the different product categories. DID identifies the average treatment effect on the treated (ATT), that is, the effect of the fat tax on the Danish consumers.
Another noteworthy advantage of our study stems from the availability of longitudinal data, which allows us to study how consumption behavior varies with the length of exposure to the tax and how the effects evolve over time, effectively embracing an event-type study approach. We adopt the estimator for event-study designs proposed by Callaway and Sant’Anna, (2021), which extends the doubly robust estimator of Sant’Anna and Zhao, (2020) to the setting with multiple time periods. [^10] Note that Callaway and Sant’Anna, (2021) additionally consider the setting with variation in treatment timing (staggered adoption), where individuals belong to a certain group depending on the timing of their treatment. However, in our context, the number of groups equals one as the tax was simultaneously introduced and repealed for all individuals. As we have already mentioned in the introduction, the main advantage of this approach over other estimators, such as the common two-way fixed effects estimator (see, for instance, Roth et al., (2023) for a recent review of the DID literature), is that it allows to control for observed covariates in a flexible way.
The quarterly panel we use ranges from 2009 to 2014, encompassing 20 quarters. Denote a particular quarter with $q$ where $q\in\{2009-Q1,\allowbreak 2009-Q2,\allowbreak 2009-Q3,\allowbreak...,\allowbreak 2014-Q4\}$. Denote with $Y_{iq}$ the outcome of interest for household $i$ in quarter $q$. As described in Section 3.1, three different outcomes concerning consumption behavior (weight, packages, and expenditure) are studied. In addition we also estimate the effect of the tax on average prices paid per 100 units (gr or ml depending on the product). The tax, our treatment, was introduced at the beginning of the fourth quarter of 2011, meaning that the third quarter of 2011 ($32011\text{-}Q$ 3) was the last pre-treatment quarter, and it was abolished at the beginning of the first quarter of 2013; thus, $42012\text{-}Q$ 4 is the last period where the tax was still in place. Let $D_{i}=1$ if unit $i$ is a Danish household, and $D_{i}=0$ otherwise. Define $Y_{pre}$ as the reference period, e.g., $32011\text{-}Q$ 3. It is important to note that the reference period varies depending on the causal effect of interest, and the precise reference period used will be explicitly stated on a case-by-case basis. Ignoring for the moment pre-treatment covariates, when all the treated units are treated at the same time, as in our case, the $ATT(q)$ are identical to the ATTs in the two-period DID approach and are identified by:
$ATT(q)=E(Y_{q}-Y_{pre}|D=1)-E(Y_{q}-Y_{pre}|D=0).$ (1)
When controlling for pre-treatment covariates, denoted by $X$, the ATT in each quarter is identified by the following expression:
$ATT(q)=E\left[\left(\frac{D}{E(D)}-\frac{\frac{p(X)(1-D)}{1-p(X)}}{E\left[\frac{p(X)(1-D)}{1-p(X)}\right]}\right)(Y_{q}-Y_{pre}-m_{q}(X))\right],$ (2)
where the propensity score $p(X)$ denotes the probability of being subject to the tax, conditional on the pre-tax households characteristics $X$, and $m_{q}(X)=E(Y_{q}-Y_{pre}|X,D=0)$ denotes the population outcome regression for the Northern-German households in quarter $q$.
Equations (1) and (2) provide the ATT for each individual quarter. A straightforward approach to establish the overall average causal effect for a specific period is to aggregate the time average treatment effects, $ATT(q)$, of that period. For example, we can aggregate the $ATT(q)$ in the tax period as follows:
$ATT=\frac{1}{5}\sum_{q\in\{2011-Q4,...,2012-Q4\}}ATT(q).$ (3)
Simply put, this is the mean over the $ATT(q)$ for the period that the tax was in effect.
Estimation proceeds in two steps. In the first step, the propensity score $p(x)$ and outcome model $m_{q}(X)$ are estimated using logistic and linear regressions, respectively. In the second step, the fitted values of the two models are plugged into the sample analogue of Equation (2). The doubly robust DID estimator attains the semiparametric efficiency bound when both nuisance models are correctly specified. A strength of using Callaway and Sant’Anna, (2021) is that they provide simultaneous confidence bands which account for multiple-testing and are obtained by a multiplier-type bootstrap procedure.
Some assumptions that are standard in literature need to be satisfied for Equations (1) to (3) to be valid. The first assumption rules out anticipation effects. This implies that Danish households did not adjust their consumption behavior in anticipation of the tax, e.g. by hoarding. As the tax was adopted approximately half a year before its implementation, we cannot exclude the presence of hoarding. We rigorously study the presence of anticipation effects of the fat tax in Section 4.1.
The second assumption is the parallel trends assumption. This assumption states that in the absence of the Danish fat tax, the outcome trends in Denmark would have followed the same trends as the ones observed in Northern-Germany during the tax period. If the composition of the treatment and control groups are different in terms of household characteristics, then this might cause the German and Danish trends to diverge in case of heterogeneous treatment effects, leading to a violation of the unconditional parallel trends assumption (see Roth et al., (2023)). To address this concern, we relax the parallel trends assumption slightly and allow for the presence of covariate-specific trends by conditioning on pre-tax household characteristics. When controlling for observed covariates we also need to impose the standard overlap assumption, i.e., it must hold that for all $X$ that $p(X)<0$. This assumption ensures that for each Danish household, we can find German households with similar characteristics. We investigate the plausibility of the parallel trends and overlap assumptions in Section 4.1.
It is crucial to note that tests for the parallel trends assumption offer insights into pre-tax trend similarities. However, the parallel trends assumption is essentially untestable as the post treatment trends that Denmark would have experienced had the tax not been introduced cannot be observed. This means that this assumption could be perfectly plausible in the pre-tax period, while violated in the post-tax period, and vice-versa. Nonetheless, observing similar pre-tax trends provides evidence in favor of the parallel trends assumption.
In Section 5 we run further robustness checks. First, we exclude Danish households living close to the German border, since the introduction of a tax can lead to increased border trade to avoid the tax. Second, we investigate how much our results rely on controlling for household characteristics by estimating the effects unconditionally.
4 Results
In this section, we report the main results. [^11] We tailored the R package of Callaway and Sant’Anna, (2021) to our setting and express our gratitude to the original authors for making their code publicly available.
4.1 Plausibility of identifying assumptions
We first examine the plausibility of the no anticipation assumption which is crucial for DID. As mentioned in Section 3.2, this assumption posits that households did not change their behavior in anticipation of the tax implementation. To explore this, we estimate the effect of the tax in the two quarters preceding its introduction, namely the second and third quarters of 2011. The tax was announced on March 17, 2011, and introduced in October, 2011. The results reported in Table 1 do not show any statistically significant anticipation effect. Nevertheless, we take a conservative stand and, to avoid any potential bias due to anticipation effects, we exclude the second and third quarter of 2011 from the analysis. We find qualitatively similar results including these two quarters. [^12] The results are available from the authors upon request.
Table 1: The causal effect of the Danish fat tax in the two quarters preceding the tax, using the first quarter of 2011 as reference period.
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Quarterly average weight consumed in gr/ml $ATT_{2011-Q2}$13.910 82.603 -142.069 -45.696 38.916 -54.424 54.246 106.351 (se) (47.829) (107.047) (160.676) (94.261) (60.690) (124.059) (47.474) (51.176) $ATT_{2011-Q3}$ 26.963 95.758 -99.659 -37.812 45.896 202.126 76.975 72.012 (se) (42.933) (100.254) (143.327) (87.195) (63.777) (135.263) (49.936) (48.075) Panel B: Quarterly average expenditure in Euro cents $ATT_{2011-Q2}$19.721 14.544 -48.137 -13.068 32.325 -44.864 71.076 28.082 (se) (39.060) (49.050) (130.123) (28.545) (38.540) (28.308) (48.302) (14.516) $ATT_{2011-Q3}$ 8.636 51.837 -12.612 -4.428 12.671 39.789 99.147 20.986 (se) (36.968) (47.847) (133.483) (28.970) (40.789) (30.300) (47.703) (13.646) Panel C: Quarterly average amount (in packages) consumed $ATT_{2011-Q2}$ 0.009 0.243 -0.145 -0.309 0.087 -0.095 0.317 0.263 (se) (0.231) (0.419) (0.497) (0.366) (0.255) (0.269) (0.296) (0.207) $ATT_{2011-Q3}$ 0.182 0.196 -0.278 -0.200 0.059 0.405 0.522 0.197 (se) (0.188) (0.427) (0.484) (0.320) (0.270) (0.283) (0.279) (0.185) Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors are clustered at household level. The results are obtained controlling for the age of the head of the household, the number of children under 15 years, the gender of diary keeper, the educational level (ISCED), the income level, and the household size.
A second aforementioned crucial assumption for the credibility of our estimates is the parallel trends assumption. An advantage of having multiple pre-tax time periods is the opportunity to gain some insight into the plausibility of this assumption. We assess the plausibility in two different ways. First, Table 2 contains the p-values of a Wald test for parallel trends, where the assumption of parallel trends in the pre-tax period is tested under the null hypothesis that they are parallel. Notably, the p-values derived from this approach do not raise concerns, as we fail to reject the null hypothesis of parallel trends, with the exception of butter in the expenditure outcome. Second, the results of pseudo treatment effects for the pre-tax periods are shown in Appendix C. Reassuringly, the uniform confidence intervals in all pre-tax periods contain zero, offering further evidence in support of the parallel trends assumption for all product categories and all outcomes. Altogether, inspection of the pre-tax trends lends strong support of the plausibility of parallel trends between Danish and Northern-German households in the tax period.
Finally, we address the overlap assumption, which is commonly tested by ensuring sufficient overlap in the treated and control units’ propensity score distributions. Histograms displaying the fitted propensity scores for the Danish and Northern-German households can be found in Appendix B. The histograms show good overlap.
In conclusion, we do not find any substantial empirical evidence against the plausibility of the assumptions underlying our DID identification strategy.
4.2 The causal effect of the fat tax on consumer responses during the tax period
The results depicted in Panel A of Table 2 reveal a statistically significant decrease in the consumption of cheese, liver sausage, and salami during the tax period. Specifically, there is an average reduction of approximately 13% for cheese, 11% for liver sausage, and 12% for salami compared to the pre-tax quarter. While we do not observe significant changes in expenditure for cheese and salami, there was a noteworthy 13.75% (about 0.8 Euro) reduction in expenditure for liver sausage, as illustrated in Table 2, Panel B. Conversely, we observe a substantial increase in expenditure for butter (2.34 Euro, 25.2%), cream (0.7 Euro, 14.49%), margarine (0.8 Euro, 15.42%), and sour cream (0.27 Euro, 16%) despite the absence of any discernible impact on the consumption of these products.
Table 2: The causal effect of the Danish fat tax on the average quarterly consumption and expenditure during the tax period (October 2011 to December 2012).
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Quarterly average weight consumed in gr/ml Danish fat tax (ATT) $-56.891$ 73.292 $-351.109^{*}$ 27.663 $-120.925^{*}$ 114.433 $-66.774^{**}$ 27.768 (se) (35.788) (78.927) (127.263) (65.761) (47.718) (101.072) (36.823) (36.351) % change -10.07% 4.81% -13.05% 2.01% -11.07% 6.24% -12.10% 4.85% obs. 29,645 40,929 32,129 39,274 31,246 38,687 31,243 36,121 p-val. pre-trend 0.273 0.384 0.344 0.455 0.846 0.772 0.963 0.208 Panel B: Quarterly average expenditure in Euro cents Danish fat tax (ATT) -25.075 $234.879^{***}$ $-240.166$ $70.823^{**}$ $-81.437^{*}$ $84.643^{***}$ $-50.452$ $27.131^{*}$ (se) (26.216) (37.992) (105.520) (19.589) (28.147) (22.041) (38.160) (10.413) % change -6.85% 25.20% -10.42% 14.49% -13.75% 15.42% -8.81% 16.03% obs. 29,645 40,929 32,129 39,274 31,246 38,687 31,243 36,121 p-val. pre-trend 0.653 0.000 0.505 0.748 0.929 0.603 0.921 0.075 Panel C: Quarterly average amount (in packages) consumed Danish fat tax (ATT) -0.010 0.361 -0.304 -0.202 $-0.463$ $0.404$ -0.009 0.059 (se) (0.160) (0.314) (0.398) (0.272) (0.199) (0.198) (0.233) (0.145) % change -0.51% 6.01% -7.09% -6.19% -14.40% 10.39% -0.33% 4.77% obs. 29,645 40,929 32,129 39,274 31,246 38,687 31,243 36,121 p-val. pre-trend 0.987 0.557 0.249 0.954 0.984 0.795 0.977 0.512 Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors clustered at unit level. % change compares the estimated tax average to the pre-tax sample mean for Danish households. The estimates are obtained using the doubly robust DID approach, controlling for age of head of household, number of children under 15, gender, ISCED, income level, and household size.
Notably, there is no observed statistically significant impact on the quantity of packages purchased across all products. Nevertheless, the observed decrease in consumption of cheese, liver sausage, and salami hints at a potential adjustment by producers who may have responded to the imposed tax by reducing the package size of these three items. In contrast, our results suggest that the package sizes for butter, cream, margarine, and sour cream remained unaltered, as there was no significant impact on either the number of packages purchased or the weight consumed. At the same time, we observe a significant rise in expenditure, possibly driven by elevated prices and a low elasticity of demand for these particular products.
In summary, our findings indicate that the tax successfully achieved the intended reduction in the consumption of cheese, liver sausage, and salami during the tax period. However, it did not yield the desired effect on the consumption of the other products under examination. Notably, the tax led to increased consumer expenditure on butter, cream, margarine, and sour cream without significantly affecting the consumption of these particular products.
4.3 The causal effect of the fat tax on consumer responses after its abolishment
Table 3: The causal effect of the Danish fat tax on the average quarterly consumption and expenditure during the two years after abolishment of the tax (January 2013 to December 2014), using the final tax period (October 2012 to December 2012) as reference period.
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Quarterly average weight consumed in gr/ml Danish fat tax (ATT) $124.442^{**}$ 5.298 $557.856^{***}$35.433 $195.790^{***}$ 72.082 $131.118^{**}$ 78.072 (se) (42.177) (86.445) (115.098) (89.103) (50.680) (98.307) (40.289) (40.092) % change 23.43% 0.33% $2$54% 19.36% 4.20% 25.39% 13.17% obs. 22,354 28,507 23,735 27,544 23,315 27,085 23,312 25,831 p-val. pre-trend 0.310 0.000 0.296 0.335 0.334 0.018 0.414 0.564 Panel B: Quarterly average expenditure in Euro cents Danish fat tax (ATT) 57.635 $-171.538^{***}$ $415.404^{***}$24.235 $122.264^{**}$ $-67.918^{**}$ $136.449^{***}$ 7.812 (se) (29.636) (42.352) (92.873) (26.516) (33.875) (22.868) (38.423) (11.110) % change 15.85% -16.77% 19.19% -4.51% 22.67% -13.20% 25.21% 4.28% obs. 22,354 28,507 23,735 27,544 23,315 27,085 23,312 25,831 p-val. pre-trend 0.456 0.000 0.358 0.833 0.051 0.013 0.252 0.093 Panel C: Quarterly average amount (in packages) consumed Danish fat tax (ATT) $0.430^{**}$0.017 $0.938^{*}$ 0.145 $0.782^{**}$ 0.144 $0.654^{**}$ 0.341 (se) (0.158) (0.340) (0.393) (0.327) (0.255) (0.187) (0.234) (0.175) % change 21.53% -0.26% 21.14% 4.47% 26.19% 3.87% 23.69% 26.58% obs. 22,354 28,507 23,735 27,544 23,315 27,085 23,312 25,831 p-val. pre-trend 0.128 0.000 0.914 0.956 0.577 0.002 0.881 0.495 Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors clustered at unit level. % change compares the estimated tax average to the pre-tax sample mean for Danish households. The estimates are obtained using the doubly robust DID approach, controlling for age of head of household, number of children under 15, gender, ISCED, income level, and household size.
Table 4: The causal effect of the Danish fat tax on the average quarterly consumption and expenditure during the two years after abolishment of the tax (January 2013 to December 2014), using the pre-tax period (January 2011 to March 2011) as reference period.
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Quarterly average weight consumed in gr/ml Danish fat tax (ATT) 0.586 $256.953^{**}$ 31.182 117.048 1.505 $325.944^{**}$ 5.839 $90.648^{*}$ (se) (35.004) (72.408) (129.746) (71.148) (46.676) (92.464) (37.856) (35.617) % change 0.10% 16.05% 1.15% 8.31% 0.14% 18.05% 1.05% 15.41% obs. 33,510 45,295 36,043 43,648 35,160 42,954 35,175 40,299 p-val. pre-trend 0.182 0.362 0.268 0.369 0.859 0.853 0.959 0.138 Panel B: Quarterly average expenditure in Euro cents Danish fat tax (ATT) 0.024 $162.860^{***}$ 12.967 $57.764^{**}$33.561 39.009 21.002 $37.947^{**}$ (se) (26.678) (38.977) (105.804) (19.152) (30.410) (21.705) (38.239) (10.451) % change 0.01% 16.92% 0.56% 11.39% -5.73% 7.66% 3.59% 21.65% obs. 33,510 45,295 36,043 43,648 35,160 42,954 35,175 40,299 p-val. pre-trend 0.516 0.000 0.235 0.668 0.950 0.734 0.905 0.050 Panel C: Quarterly average amount (in packages) consumed Danish fat tax (ATT) 0.175 $1.348^{***}$ 0.586 0.058 -0.001 $0.916^{***}$ $0.597^{*}$ 0.285 (se) (0.150) (0.294) (0.368) (0.254) (0.196) (0.210) (0.222) (0.145) % change 8.53% 20.86% 12.83% 1.74% -0.04% 23.65% 21.17% 22.28% obs. 33,510 45,295 36,043 43,648 35,160 42,954 35,175 40,299 p-val. pre-trend 0.982 0.472 0.282 0.939 0.991 0.797 0.989 0.481 Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors clustered at unit level. % change compares the estimated tax average to the pre-tax sample mean for Danish households. The estimates are obtained using the doubly robust DID approach, controlling for age of head of household, number of children under 15, gender, ISCED, income level, and household size.
The fat tax was abolished in January 2013. We study the two years after the abolishment to gain insight into potential habit formation induced by the tax. We present the effects of the abolishment of the fat tax on consumption in comparison to the final quarter of the tax period in Table 3 and in comparison to the pre-tax period in Table 4.
For cheese, liver sausage, and salami, we find that consumption increased after the tax was abolished, as shown in Panel A of Table 3. This increase brought the consumption of these products back to pre-tax levels, as indicated in Panel A of Table 4. Regarding salami, we observe that the abolishment of the tax resulted in a consistent reduction in packaging size, with a 23.69% increase in the average number of packages purchased. However, this change had no significant effect on the average consumption in terms of weight. Overall, despite the short-term reduction in consumption observed during the tax period, these effects seemed to have disappeared after the tax was abolished. This is in line with Allais et al., (2010), who found that long-term effects can take seven years to establish, as mentioned in the literature review.
Panel A of Table 4 shows that butter, margarine and sour cream experienced a substantial increase in consumption after the abolishment of the tax in comparison to the pre-tax period as well as an overall increase in expenditure and average number of packages purchased (see Panel B and C).
In summary, the tax failed in creating any long-term reductions in consumption of the targeted products, and its abolishment induced a short-term increase in consumption for some of these products.
4.4 Price mechanism
Since the tax was imposed on producers, it remains unclear to what extent consumers experienced price increases for the taxed products. In this section, we aim to estimate the causal effect of the tax on the prices paid for each product. One limitation of this approach is that we can only observe prices for products that have been purchased, not the overall price level. However, the results presented in Table 5 provide us with insights into whether consumers encountered higher prices for the products they chose to purchase.
Table 5 shows that prices for butter, cream, margarine, and sour cream significantly increased in response to the tax by 23.44%, 10.58% , 11.35%, and 12.31%, respectively. These price increases are lower than those anticipated by the Danish Tax Committee, who anticipated increases in consumer prices up to 35% for animal fat, 33% for dairy, and 24% for margarine. The price increases to a large extent explain our previous finding that consumer expenditure on these products increased by 16%, 22%, and 10%, respectively. The rise in prices corresponded with a notable expenditure increase for both butter and margarine products. Evidently, consumers were willing to pay higher prices for these particular products. The price increase persisted also after the tax was abolished for butter and sour cream, while we find an overall decrease for margarine. We do not find a significant increase in price for other products during the tax period. Interestingly, we observe a substantial decrease in liver sausage prices of about 18%.
Table 5: The causal effect of the Danish fat tax on the quarterly average price paid per 100 units.
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Tax period versus pre-tax quarter Danish fat tax (ATT) -1.918 $14.935^{***}$ $3.610$ $3.968^{***}$3.076 $3.566^{***}$ 1.976 $3.863^{***}$ (se) (2.916) (0.558) (1.660) (0.408) (2.478) (0.509) (3.292) (0.586) % change -2.81% 23.44% 3.92% 10.58% -4.95% 11.35% 1.65% 12.31% obs. 16,863 32,189 27,511 27,735 22,726 28,682 20,925 19,590 p-val. pre-trend 0.554 0.000 0.683 0.009 0.718 0.441 0.833 0.000 Panel B: Post-tax period versus last taxed quarter Danish fat tax (ATT) -3.444 $-10.671^{***}$2.434 $-1.701^{***}$3.917 $-4.054^{***}$3.465 $-1.898^{***}$ (se) (2.894) (0.676) (1.996) (0.474) (2.716) (0.569) (3.595) (0.530) % change -4.82% -15.79% -2.58% -4.18% -6.27% -13.03% -2.92% -5.83% obs. 12,568 22,942 19,930 19,452 16,170 19,230 15,208 13,678 p-val. pre-trend 0.787 0.000 0.489 0.000 0.538 0.032 0.533 0.000 Panel C: Post-tax period versus pre-tax quarter Danish fat tax (ATT) -6.734 $2.039^{**}$ 0.086 -0.093 $-11.091^{***}$ $-1.483^{**}$5.631 $2.585^{***}$ (se) (2.889) (0.583) (1.644) (0.371) (2.414) (0.484) (3.351) (0.572) % change -9.67% 3.22% 0.09% -0.24% -17.95% -5.01% -4.73% 8.15% obs. 19,235 36,058 31,104 30,979 25,609 31,466 23,729 21,739 p-val. pre-trend 0.723 0.000 0.757 0.009 0.615 0.557 0.820 0.000 Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors clustered at unit level. % change compares the estimated tax average to the pre-tax sample mean for Danish households. The estimates are obtained using the doubly robust DID approach, controlling for age of head of household, number of children under 15, gender, ISCED, income level, and household size.
Overall, we do not find a significant increase in prices for products where we observed a short run decrease in consumption. Since we use prices of products purchased by households, the effect on prices might actually be downward biased, because households might decide not to buy when confronted with excessively high prices.
5 Robustness checks
In this section, we conduct several robustness checks to assess the reliability of the main findings presented in the previous section, addressing the potential concerns with respect to our identification strategy stated in Section 3.2. First, the issue of potential cross-border shopping merits attention, as it might obscure the effect of the tax. To investigate potential cross-border shopping effects, we exclude Southern-Danish households from South/West Jutland and Funen, focusing on Danish households that are less likely to engage in cross-border shopping and hence fully subject to the tax. Interestingly, the results in Table 11 in Appendix D.1 exhibit similar patterns as our main results, indicating that changes in consumption behavior were similar for Danish households that were able to cross-border shop compared to Danish households that did not have this opportunity. Note that in our main results, the unintended effect of cross-border shopping is incorporated as we do include purchases made abroad.
Second, we consider concerns surrounding disparities in Danish and Northern-German household characteristics and assess whether differences in household characteristics between Danish and Northern-German households produced the effects of our main results in Section 4. To gauge the influence of the covariates, we contrast the treatment effects obtained under the conditional parallel trends assumption with those derived under the unconditional parallel trends assumption. If doubly robust DID estimates based on the conditional parallel trends assumption resemble the estimates based on the unconditional parallel trends assumption, this would indicate that covariates contain limited information on consumer behavior trends and are unlikely to drive the results in Section 4. For brevity, we only included the results for the weight outcome, but similar observations apply to the other two main outcomes. Interested readers may obtain the complete set of results upon request. Reassuringly, Figure 39 in Appendix D.2 display extremely similar trends compared to the figures in Appendix C. This suggests that covariates have minimal influence on the trends. Upon comparing the results, it becomes evident that, unsurprisingly, incorporating covariates enhances precision.
6 Heterogeneity analysis by income levels
Criticisms of the tax being regressive suggest that low-income households may be disproportionately affected compared to wealthier ones. Groceries constitute a larger proportion of the income of poorer households, making them disproportionately affected by the tax if they do not adjust their consumption behavior (Allais et al., , 2010; Zhen et al., , 2011). High-income households, on the other hand, might respond more to health incentives rather than monetary ones. Understanding how the tax affects different income groups is vital for targeted policy improvements, and extrapolating the tax’s applicability to other countries.
(a) Average quarterly weight consumed (gr/ml).
(b) Average quarterly expenditure in Euro cents.
(c) Average quarterly number of packages.
Figure 10: Doubly robust DID estimates of the causal effect of the Danish fat tax during the tax period (October 2011 to December 2012) for low, medium and high-income households. Low-income households are defined as households with very low income or low income. High-income households are defined as households with very high income or high income. 95% confidence intervals in red.
(a) Average quarterly weight consumed (gr/ml).
(b) Average quarterly expenditure in Euro cents.
(c) Average quarterly number of packages.
Figure 11: Doubly robust DID estimates of the causal effect of the Danish fat tax during the post-tax period (January 2013 to December 2014) compared to the pre-tax quarter (January 2011 to March 2011) for low, medium and high-income households. Low-income households are defined as households with very low income or low income. High-income households are defined as households with very high income or high income. 95% confidence intervals in red.
In order to investigate potential heterogeneity in the effects of the Danish fat tax, we estimate the effect in two different subsets of the data. First, we estimate the effects of the tax on our three consumption outcomes in the subsample of households who have a low or very low income (refer to Section 3.1 to see how we define income categories). Second, we estimate the same effects but for households who have high or very-high income. The results are presented in Figure 10. Interestingly, Panel (a) of Figure 10 shows a significant reduction in bacon consumption for low-income households, for whom we also observe a significant decrease in consumption of cheese and liver sausage. The effects are not statistically significant for these products, for neither medium-income nor high-income households. Moreover, we find that low-income and medium-income households also reduced their expenditure on cheese and liver sausage, as shown in Panel (b) of Figure 10.
Similar to the main analysis, we observe an uptick in expenditure for both butter and margarine. Notably, the increase is fairly similar between households with high and low incomes, but not statistically significant for medium-income households. This means that low-income households were disproportionately affected by the tax.
The results in Figure 11 show that, after the abolishment of the tax, the increase in butter, cream, margarine, and sour cream consumption is largely driven by high-income households, with relatively weaker effects on low-income households. Interestingly, we find a decrease in the consumption of bacon only for low-income households. However, the effect is marginally significant. [^13] For the interested reader, Figure 40 in Appendix E provides the heterogeneity results for the abolishment effects when comparing the post-tax period to the last taxed quarter.
7 Conclusion
The present study has been the first to estimate the causal effect of the Danish tax targeting saturated fats on consumer behavior and expenditure, employing a valid and culturally-close geographic control group consisting of Northern-German households, located in Schleswig-Holstein and Hamburg. Our analysis is based on Callaway and Sant’Anna, ’s (2021) doubly robust DID estimators that allow to flexibly include observed covariates and provide confidence intervals that account for multiple hypothesis testing. The longitudinal nature of our data allowed us to gain valuable insights into the evolution of causal effects over time. To ensure the validity of our findings, we studied the plausibility of the identifying assumptions, most crucially the parallel trends and no anticipation assumption, and performed robustness checks with respect to cross-border shopping.
Distinguishing itself as the first to investigate expenditure patterns, tax abolishment effects, and sustained habit formation post-tax, our research contributes a comprehensive analysis of the policy’s implications. Findings on the purchased weight, the number of packages purchased, and the expenditure paint an interesting picture of the effect of the tax on consumer behavior. Perhaps the most relevant finding is that the tax achieved a reduction in consumption of bacon, cheese, and liver sausage, and salami ranging between 11% and 13%. For bacon, cheese, and liver sausage, the effect of the tax was mainly driven by low-income consumers. This indicates that the tax aligned with its goal of reducing the intake of saturated fats from these products in the population that might benefit most from a reduction. It is important to emphasize, however, that it is unclear whether these products are substituted by healthier alternatives, offering an important avenue for future research. After the tax was repealed, these effects are no longer statistically significant.
The second most significant finding to emerge from this study is the unchanged consumption of margarine, butter, sour cream, and cream, while expenditure on these products increased between 14% and 25%. Interestingly, the increase in expenditure for these products is similar for low and higher-income households, which implies the tax’s regressivity with respect to these products. There are various reasons for this observation, and to improve the design of fat taxes in future, more research is needed to understand why low-income consumers do not substitute away from these products. Exploring the mechanisms behind observed consumer responses is a promising avenue for future research. For instance, gaining information on final consumer characteristics, rather than household characteristics, could provide deeper insights.
Previous studies exclusively focused on the tax’s implementation phase, neglecting post-tax dynamics. Our study, however, sheds light on intriguing post-tax patterns. Notably, the removal of the tax led to varied effects across different product categories. An important finding showed elevated quantities consumed of products high in saturated fat content, even surpassing pre-tax levels in some cases. This raises concerns that the process of implementing and abolishing the tax might have exacerbated unfavorable consumer behavior. For instance, we did observe persisting elevated consumption patterns for butter, margarine, and sour cream in the two years post-tax, indicating that there might be some tax-induced habit formation with respect to this product, especially for high-income households. In future research, it could be interesting to further investigate the longer-term impact (i.e., the impact beyond the year 2014) of the tax’s abolishment.
Our examination of price dynamics has revealed that not all price changes are passed on to the consumers and reflected in the store prices. Among certain product categories (butter, cream, margarine, and sour cream), we noted price increases, while other products displayed no significant price alterations. The abolishment of the tax again resulted in divergent price changes. While prices decreased for some products, such as margarine and liver sausage, others, like butter and sour cream, exhibited persistent elevated prices.
Understanding these effects and the underlying mechanisms is imperative for policymakers. A food tax can be used as a tool to improve public health by influencing consumers’ behavior while increasing revenue from tax. Taken together, our findings do not support strong recommendations to impose a broad tax on saturated fat, as the patterns in consumer responses vary heavily across products. To optimize the design of a tax on saturated fats, more research is required to understand the mechanisms driving consumer choices and whether substitutions towards healthier alternatives are made. Equally important is the consideration of the policy’s impact on low-income households, as our results clearly show that they are disproportionately affected by the tax.
At present, several questions persist regarding the impact of the fat tax. While our analysis considered specific aggregated food categories, an intriguing avenue for future research involves studying substitution dynamics within these product categories. Unfortunately, the study’s scope was constrained by the lack of detailed data on the fat percentages of products within these categories. Access to data on fat percentages could also enable a direct comparison between observed price changes and anticipated adjustments if the tax were entirely shifted to consumers, providing a more thorough analysis of the tax’s pricing dynamics. Furthermore, more research is required to determine the implications of the tax on health outcomes. An extended investigation into health-related consequences, such as potential effects on BMI, would offer a more comprehensive understanding of the tax’s potential long-term health implications. Finally, understanding the impact of health messaging induced by the fat tax and how the public debate shaped attitudes toward products high in saturated fat is a crucial aspect that warrants further study.
Appendix A Data processing steps and descriptives
A.1 Data processing steps
Figure 12: Data flow chart with the number of observations per step for constructing the Danish data set covering January 2009 to December 2014. For each product category, a quarterly panel is created which includes households that purchase some product from that category at least once between January 2009 and December 2014.
some text such that figure appears at top of page
Figure 13: Data flow chart with the number of observations per step for constructing the Northern-German data set covering January 2009 to December 2014. For each product category, a quarterly panel is created which includes households that purchase some product from that category at least once between January 2009 and December 2014.
A.2 Summary statistics of pre-tax household characteristics
Table 6: Summary statistics of pre-tax socio-economic household characteristics for Danish and Northern-German households used in the main analysis in Table 2.
Variable Levels $\mathbf{n_{\mathrm{Denmark}}}$$\mathbf{\%_{\mathrm{Denmark}}}$ $\mathbf{n_{\mathrm{Germany}}}$$\mathbf{\%_{\mathrm{Germany}}}$ $\mathbf{n_{\mathrm{all}}}$$\mathbf{\%_{\mathrm{all}}}$
Geographic region Capital region 575 21.3 0 0.0 575 12.8
Zealand 606 22.5 0 0.0 606 13.5
Funen 276 10.2 0 0.0 276 6.2
North Jutland 382 14.2 0 0.0 382 8.5
East Jutland 456 16.9 0 0.0 456 10.2
South/West Jutland 402 14.9 0 0.0 402 9.0
Hamburg 0 0.0 661 37.0 661 14.8
Schleswig-Holstein 0 0.0 1,123 63.0 1,123 25.1
Age (head of hh.) Under 30 573 21.2 202 11.3 775 17.3
30-39 713 26.4 362 20.3 1,075 24.0
40-49 609 22.6 468 26.2 1,077 24.0
50-59 512 19.0 333 18.7 845 18.9
60-69 136 5.0 304 17.0 440 9.8
70 or over 154 5.7 115 6.4 269 6.0
Children under 15 1 2,131 79.0 1,300 72.9 3,431 76.6
2 253 9.4 280 15.7 533 11.9
3 224 8.3 154 8.6 378 8.4
4 82 3.0 45 2.5 127 2.8
5 7 0.3 4 0.2 11 0.2
7 0 0.0 1 0.1 1 0.0
Gender (diary keeper) Male 599 22.2 737 41.3 1,336 29.8
Female 2,098 77.8 1,047 58.7 3,145 70.2
Education 2 543 20.1 413 23.1 956 21.3
(ISCED level) 3 1,011 37.5 666 37.3 1,677 37.4
5 392 14.5 175 9.8 567 12.6
6 578 21.4 211 11.8 789 17.6
7 173 6.4 319 17.9 492 11.0
Income (head of hh.) Very low income 837 31.0 608 34.1 1,445 32.2
Low income 547 20.3 344 19.3 891 19.9
Medium income 361 13.4 298 16.7 659 14.7
High income 451 16.7 214 12.0 665 14.8
Very high income 501 18.6 320 17.9 821 18.3
Household size 1 981 36.4 455 25.5 1,436 32.0
2 1,045 38.8 689 38.6 1,734 38.7
3 274 10.2 327 18.3 601 13.4
4 286 10.6 226 12.7 512 11.4
5 92 3.4 62 3.5 154 3.4
6 18 0.7 20 1.1 38 0.8
7 1 0.0 4 0.2 5 0.1
8 0 0.0 1 0.1 1 0.0
Country Denmark 2,697 100.0 0 0.0 2,697 60.2
Germany 0 0.0 1,784 100.0 1,784 39.8
all 2,697 100.0 1,784 100.0 4,481 100.0
A.3 Summary statistics for the consumption and price outcomes during the pre-tax and tax period
Table 7: This table contains the average total quarterly consumption per household measured by the average weight and the average amount (counted by number of packages) purchased before and after the tax introduction. It further contains the average quarterly expenditure in Euro cents per household and the average price paid per 100 units. The pre-tax period covers January 2009 to September 2011, and the tax period covers October 2011 to December 2012.
Product Outcome Pre-tax DK Tax DK $\Delta_{\text{DK}}$ Pre-tax DE Tax DE $\Delta_{\text{DE}}$ Bacon Weight 562.67 486.51 -76.16 382.82 371.16 -11.66 Amount 1.89 1.80 -0.09 1.71 1.68 -0.03 Expenditure 359.84 331.98 -27.86 268.52 274.60 6.08 Price 67.74 71 3.26 79.93 85.08 5.15 Butter Weight 1,607.62 1,554.70 -52.92 1,507.32 1,483.01 -24.31 Amount 6.27 6.22 -0.05 6.13 6.00 -0.13 Expenditure 913.97 1,068.94 154.97 595.71 625.36 29.65 Price 59.65 72.36 12.71 41.06 43.41 2.35 Cheese Weight 2,717.23 2,146.68 -570.55 2,678.16 2,569.77 -108.39 Amount 4.24 3.76 -0.48 9.44 9.30 -0.14 Expenditure 2,306.91 1,912.95 -393.96 1,942.41 1,948.03 5.62 Price 90.50 95.79 5.29 74.37 78.13 3.76 Cream Weight 1,465.81 1,548.64 82.83 1,099.53 1,108.92 9.39 Amount 3.47 3.53 0.06 5.10 5.14 0.04 Expenditure 481.91 569.35 87.44 242.29 267.57 25.28 Price 34.95 39.24 4.29 23.03 25.14 2.11 Liver sausage Weight 1,138.17 977.71 -160.46 538.43 549.05 10.62 Amount 3.31 2.84 -0.47 2.87 2.89 0.02 Expenditure 613.31 530.67 -82.64 374.37 402.22 27.85 Price 61.47 63.19 1.72 75.21 79.22 4.01 Margarine Weight 1,842.00 1,693.05 -148.95 1,969.10 1,760.18 -208.92 Amount 3.87 3.63 -0.24 4.12 3.58 -0.54 Expenditure 511.17 570.55 59.38 368.39 350.34 -18.05 Price 29.35 34.78 5.43 19.87 21.29 1.42 Salami Weight 568.87 480.48 -88.39 490.46 477.05 -13.41 Amount 2.69 2.47 -0.22 3.12 3.02 -0.10 Expenditure 598.54 506.25 -92.29 465.39 466.93 1.54 Price 121.07 121.47 0.40 111.13 113.61 2.48 Sour cream Weight 613.76 604.32 -9.44 506.46 524.83 18.37 Amount 1.33 1.30 -0.03 2.47 2.55 0.08 Expenditure 176.17 186.28 10.11 127.57 132.27 4.70 Price 30.57 32.43 1.86 26.01 25.96 -0.05 DK indicates Danish households and DE indicates Northern-German households. $\Delta$ indicates the difference.
Appendix B Fitted propensity scores for overlap verification
Figure 14: Bacon.
Figure 15: Butter.
Figure 16: Cheese.
Figure 17: Cream.
Figure 18: Liver sausage.
Figure 19: Margarine.
Figure 20: Salami.
Figure 14: Bacon.
Figure 15: Butter.
Figure 16: Cheese.
Figure 17: Cream.
Figure 18: Liver sausage.
Figure 19: Margarine.
Figure 20: Salami.
Figure 21: Sour cream.
Figure 22: Fitted propensity scores for the Danish and Northern-German households obtained with a logistic regression controlling for the age of the head of the household, the number of children under 15, gender, educational attainment (ISCED), income level, and household size. The overlap assumption requires a positive probability of being treated at every covariate value, which seems to be valid for the cases depicted here.
Appendix C Evolution of tax effects over time
C.1 Bacon
(a) Average quarterly weight consumed (gr/ml) per household.
(b) Average quarterly expenditure in Euro cents per household.
(c) Average quarterly number of packages per household.
Figure 23: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (1) of Tables 8 to 10.
C.2 Butter
(a) Average quarterly weight consumed (gr/ml) per household.
(b) Average quarterly expenditure in Euro cents per household.
(c) Average quarterly number of packages per household.
Figure 24: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (2) of Tables 8 to 10.
C.3 Cheese
(a) Average quarterly weight consumed (gr/ml) per household.
(b) Average quarterly expenditure in Euro cents per household.
(c) Average quarterly number of packages per household.
Figure 25: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (3) of Tables 8 to 10.
C.4 Cream
(a) Average quarterly weight consumed (gr/ml) per household.
(b) Average quarterly expenditure in Euro cents per household.
(c) Average quarterly number of packages per household.
Figure 26: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (4) of Tables 8 to 10.
C.5 Liver sausage
(a) Average quarterly weight consumed (gr/ml) per household.
(b) Average quarterly expenditure in Euro cents per household.
(c) Average quarterly number of packages per household.
Figure 27: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (5) of Tables 8 to 10.
C.6 Margarine
(a) Average quarterly weight consumed (gr/ml) per household.
(b) Average quarterly expenditure in Euro cents per household.
(c) Average quarterly number of packages per household.
Figure 28: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (6) of Tables 8 to 10.
C.7 Salami
(a) Average quarterly weight consumed (gr/ml) per household.
(b) Average quarterly expenditure in Euro cents per household.
(c) Average quarterly number of packages per household.
Figure 29: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (7) of Tables 8 to 10.
C.8 Sour cream
(a) Average quarterly weight consumed (gr/ml) per household.
(b) Average quarterly expenditure in Euro cents per household.
(c) Average quarterly number of packages per household.
Figure 30: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (8) of Tables 8 to 10.
Table 8: Dynamic treatment effects, $ATT(q)$, for quarterly average weight consumed in gr/ml.
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream $ATT_{2009-Q2}$69.13 -62.88 46.73 82.69 18.71 -155.22 -40.79 108.83 (se) (49.88) (108.28) (176.68) (99.88) (65.42) (132.20) (59.06) (54.68) $ATT_{2009-Q3}$ 54.29 84.87 183.78 -61.67 30.27 42.89 44.76 -20.94 (se) (52.08) (107.77) (167.77) (98.53) (64.20) (130.56) (48.70) (53.75) $ATT_{2009-Q4}$22.12 -26.10 -205.94 141.64 -45.28 54.04 -39.73 -18.69 (se) (51.48) (95.67) (174.33) (98.05) (66.10) (127.49) (48.99) (47.99) $ATT_{2010-Q1}$ 60.76 -16.14 207.32 -13.82 19.83 39.28 25.31 52.41 (se) (55.35) (102.61) (163.09) (91.72) (65.27) (120.48) (48.77) (46.31) $ATT_{2010-Q2}$51.84 -39.90 -53.03 -120.96 -38.56 66.58 -11.71 -21.71 (se) (49.27) (93.27) (159.11) (87.73) (62.45) (125.17) (45.25) (43.43) $ATT_{2010-Q3}$79.11 75.50 -183.07 31.91 -2.71 -157.30 -32.46 -5.21 (se) (47.49) (98.59) (162.64) (93.02) (59.56) (119.92) (50.48) (44.93) $ATT_{2010-Q4}$ 98.44 -91.26 -102.94 79.42 45.10 115.30 50.27 -64.08 (se) (51.84) (109.75) (159.96) (98.80) (58.94) (129.49) (51.06) (45.00) $ATT_{2011-Q1}$51.77 -154.55 53.94 -146.26 -93.69 -131.65 -23.13 7.46 (se) (49.12) (111.97) (164.07) (89.22) (65.03) (141.44) (47.87) (52.78) $ATT_{2011-Q4}$38.20 $-268.51^{*}$166.35 9.60 -69.14 -187.42 -50.96 -6.96 (se) (47.50) (104.34) (168.12) (87.73) (62.90) (133.87) (51.35) (48.24) $ATT_{2012-Q1}$37.38 173.17 -374.49 -71.29 -118.45 176.18 -78.86 39.55 (se) (43.02) (99.00) (163.40) (87.94) (62.85) (119.92) (51.04) (47.33) $ATT_{2012-Q2}$10.37 183.92 -305.58 49.40 $-151.23^{*}$ 263.02 -10.87 25.81 (se) (42.52) (104.80) (170.58) (98.59) (59.75) (114.91) (48.00) (51.86) $ATT_{2012-Q3}$67.53 84.42 $-432.40^{*}$ 14.28 -72.74 121.99 -61.36 76.91 (se) (47.41) (97.27) (151.62) (93.62) (60.68) (126.53) (48.25) (53.78) $ATT_{2012-Q4}$ $-130.98^{*}$ 193.46 $-476.73^{*}$ 136.32 $-193.07^{*}$ 198.39 -131.82 3.52 (se) (50.01) (101.65) (161.99) (105.67) (64.65) (125.44) (53.77) (56.03) $ATT_{2013-Q1}$35.86 203.63 -190.38 14.09 -117.06 138.68 -66.08 29.40 (se) (46.82) (113.12) (172.27) (113.39) (61.08) (127.11) (52.69) (47.13) $ATT_{2013-Q2}$58.44 185.93 -252.18 4.00 -74.42 235.68 -74.10 4.28 (se) (44.69) (102.01) (175.48) (109.42) (64.31) (128.40) (55.48) (49.50) $ATT_{2013-Q3}$68.15 83.94 -240.44 99.26 -11.90 $321.98^{*}$32.42 $141.56^{*}$ (se) (47.15) (107.25) (166.99) (109.50) (65.00) (122.13) (47.74) (53.04) $ATT_{2013-Q4}$4.98 $334.66^{**}$84.59 238.34 -107.59 $408.00^{**}$69.05 70.68 (se) (53.13) (122.59) (153.75) (108.01) (66.53) (132.07) (53.49) (52.88) $ATT_{2014-Q1}$ 35.33 171.35 409.49 103.39 $199.76^{*}$ $347.68^{*}$ 66.88 60.76 (se) (50.44) (115.20) (177.38) (92.15) (68.70) (134.56) (54.76) (50.25) $ATT_{2014-Q2}$ 66.08 $373.91^{**}$ $532.98^{**}$ 235.67 $177.27^{*}$ $382.61^{*}$ 71.63 138.02 (se) (49.66) (107.07) (175.17) (100.61) (69.07) (126.95v (56.85) (57.16) $ATT_{2014-Q3}$ 55.80 239.85 $571.49^{**}$ 114.29 91.23 270.31 108.76 132.13 (se) (49.77) (106.39) (157.07) (97.36) (66.37) (136.47) (47.13) (54.38) $ATT_{2014-Q4}$ 94.84 $406.58^{**}$ 310.35 193.80 90.84 $420.97^{**}$ 61.84 34.57 (se) (54.20) (115.83) (185.60) (100.97) (68.32) (134.54) (54.52) (49.74)
Table 9: Dynamic treatment effects, $ATT(q)$, for quarterly average expenditure in Euro cents.
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream $ATT_{2009-Q2}$31.04 -21.88 -30.48 12.80 -1.18 -46.41 -33.34 33.05 (se) (36.93) (43.08) (146.61) (26.52) (42.08) (32.55) (55.09) (14.97) $ATT_{2009-Q3}$ 35.68 -11.59 84.82 -23.96 31.27 0.78 28.94 -3.56 (se) (36.72) (40.63) (131.02) (25.57) (38.95) (30.84) (49.14) (13.45) $ATT_{2009-Q4}$9.44 $-107.16^{*}$120.58 8.55 -44.16 13.18 -24.89 -9.83 (se) (36.81) (40.67) (124.97) (25.30) (38.11) (28.32) (51.16) (11.70) $ATT_{2010-Q1}$ 23.51 -0.60 144.59 7.57 20.36 -2.28 21.20 4.65 (se) (35.88) (45.53) (122.55) (27.11) (43.61) (28.83) (50.98) (12.32) $ATT_{2010-Q2}$32.24 -36.30 -59.73 -25.60 -9.85 31.54 -32.31 -4.75 (se) (35.10) (41.60) (122.78) (27.02) (40.23) (28.84) (45.35) (12.76) $ATT_{2010-Q3}$39.12 30.61 -87.50 14.24 11.29 -39.83 -24.43 -2.91 (se) (31.80) (43.56) (134.71) (26.56) (38.32) (27.40) (47.54) (13.14) $ATT_{2010-Q4}$ 48.00 22.71 -121.89 19.77 16.34 35.52 38.51 -17.71 (se) (37.24) (49.20) (135.13) (25.68) (40.65) (27.97) (51.33) (12.49) $ATT_{2011-Q1}$23.38 $-142.21^{*}$ 1.83 -44.53 -49.93 -28.71 -33.84 1.34 (se) (35.96) (51.37) (142.75) (28.18) (43.26) (31.01) (48.12) (13.06) $ATT_{2011-Q4}$25.36 5.41 -83.51 67.51 -25.99 11.52 -16.28 6.71 (se) (37.24) (50.84) (136.30) (26.61) (38.77) (30.44) (51.35) (14.04) $ATT_{2012-Q1}$14.43 $217.48^{***}$264.10 47.97 -87.25 $99.13^{**}$71.69 22.93 (se) (31.82) (48.47) (128.00) (28.60) (39.82) (28.47) (47.25) (13.76) $ATT_{2012-Q2}$ 12.74 $324.86^{***}$248.37 $86.44^{*}$ $-98.86^{*}$ $117.37^{***}$2.84 31.84 (se) (29.84) (50.69) (130.35) (29.09) (38.18) (28.61) (46.73) (13.39) $ATT_{2012-Q3}$38.24 $288.38^{***}$254.46 65.20 -45.69 $89.35^{**}$42.75 $46.27^{**}$ (se) (34.55) (49.14) (127.70) (25.62) (40.62) (29.98) (50.34) (13.97) $ATT_{2012-Q4}$60.10 $338.27^{***}$ $-350.39^{*}$ $87.00^{*}$ $-149.40^{**}$ $105.84^{**}$118.69 27.90 (se) (37.80) (56.32) (136.13) (30.84) (43.71) (28.86) (51.54) (14.39) $ATT_{2013-Q1}$32.24 $163.85^{**}$179.63 32.94 $-101.29^{*}$ 28.18 -58.35 26.79 (se) (35.70) (53.27) (142.15) (29.79) (38.01) (29.36) (57.03) (13.87) $ATT_{2013-Q2}$27.56 122.16 -219.92 16.71 -79.78 46.91 -57.49 34.10 (se) (32.52) (55.70) (134.93) (30.63) (41.05) (28.91) (48.70) (13.90) $ATT_{2013-Q3}$50.83 97.25 -145.88 56.58 -52.54 42.48 -18.54 $57.03^{***}$ (se) (36.92) (55.79) (139.93) (32.66) (36.21) (29.52) (49.46) (14.24) $ATT_{2013-Q4}$19.64 125.40 -210.05 45.55 $-108.22^{*}$ 52.37 -61.48 31.53 (se) (35.77) (58.63) (136.76) (30.07) (42.44) (28.48) (51.21) (14.14) $ATT_{2014-Q1}$ 27.81 51.14 329.05 69.48 54.22 32.90 87.94 22.87 (se) (37.11) (57.51) (144.13) (29.96) (42.00) (25.59) (52.09) (15.57) $ATT_{2014-Q2}$ 52.16 $247.94^{***}$ $506.17^{**}$ $117.15^{**}$ 77.91 45.32 102.81 $52.46^{**}$ (se) (37.18) (53.58) (147.39) (31.26) (38.47) (29.46) (51.35) (16.24) $ATT_{2014-Q3}$ 44.99 $233.53^{***}$ $480.43^{**}$ 63.47 73.67 17.97 $137.29^{*}$ $42.90^{*}$ (se) (34.80) (52.20) (139.03) (30.35) (42.48) (29.35) (50.23) (15.62) $ATT_{2014-Q4}$ 63.37 $295.09^{***}$ 297.82 61.34 34.85 38.80 84.19 21.29 (se) (39.28) (55.95) (138.15) (31.53) (40.90) (28.42) (51.81) (13.30)
Table 10: Dynamic treatment effects, $ATT(q)$, for quarterly average amount (in packages) consumed.
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream $ATT_{2009-Q2}$0.21 -0.28 -0.03 0.08 -0.02 -0.36 -0.33 0.18 (se) (0.26) (0.46) (0.52) (0.40) (0.28) (0.27) (0.31) (0.21) $ATT_{2009-Q3}$ 0.15 0.37 0.66 -0.11 0.29 0.14 0.21 0.01 (se) (0.23) (0.40) (0.51) (0.38) (0.27) (0.27) (0.29) (0.19) $ATT_{2009-Q4}$0.11 -0.12 -0.30 0.37 -0.23 0.07 -0.08 -0.05 (se) (0.25) (0.38) (0.48) (0.35) (0.27) (0.24) (0.27) (0.19) $ATT_{2010-Q1}$ 0.16 -0.03 0.79 0.01 0.05 0.07 -0.02 0.23 (se) (0.27) (0.41) (0.46) (0.35) (0.27) (0.26) (0.28) (0.18) $ATT_{2010-Q2}$0.11 -0.09 -0.08 -0.18 0.05 0.14 0.01 -0.04 (se) (0.25) (0.36) (0.48) (0.35) (0.26) (0.26) (0.28) (0.18) $ATT_{2010-Q3}$0.06 0.33 -0.39 0.18 -0.04 -0.27 -0.12 -0.05 (se) (0.22) (0.36) (0.44) (0.34) (0.27) (0.25) (0.24) (0.19) $ATT_{2010-Q4}$ 0.13 -0.38 -0.19 -0.03 0.13 0.21 0.23 -0.23 (se) (0.20) (0.41) (0.47) (0.33) (0.29) (0.26) (0.26) (0.19) $ATT_{2011-Q1}$0.11 -0.54 0.25 -0.16 -0.10 -0.31 -0.11 0.02 (se) (0.23) (0.42) (0.51) (0.35) (0.29) (0.28) (0.27) (0.18) $ATT_{2011-Q4}$ 0.07 -1.06 -0.56 -0.23 -0.33 -0.35 -0.02 -0.06 (se) (0.24) (0.44) (0.58) (0.32) (0.27) (0.28) (0.32) (0.19) $ATT_{2012-Q1}$ 0.22 0.61 0.10 -0.24 -0.40 0.50 -0.12 0.21 (se) (0.19) (0.39) (0.50) (0.33) (0.28) (0.26) (0.30) (0.19) $ATT_{2012-Q2}$ 0.07 0.70 -0.07 -0.33 -0.50 $0.74^{*}$ 0.17 0.01 (se) (0.20) (0.38) (0.47) (0.37) (0.27) (0.27) (0.30) (0.21) $ATT_{2012-Q3}$0.12 0.41 -0.44 -0.10 -0.26 0.51 0.03 0.21 (se) (0.19) (0.41) (0.47) (0.36) (0.27) (0.28) (0.33) (0.20) $ATT_{2012-Q4}$0.29 $1.14^{*}$0.55 -0.11 $-0.82^{*}$ 0.62 -0.10 -0.08 (se) (0.21) (0.42) (0.50) (0.37) (0.30) (0.26) (0.30) (0.23) $ATT_{2013-Q1}$ 0.02 1.06 0.79 -0.43 -0.30 0.60 0.15 0.04 (se) (0.22) (0.48) (0.92) (0.44) (0.29) (0.26) (0.30) (0.20) $ATT_{2013-Q2}$0.10 $1.14^{*}$0.19 -0.12 -0.20 $0.81^{**}$ 0.22 0.29 (se) (0.20) (0.39) (0.49) (0.46) (0.29) (0.26) (0.31) (0.19) $ATT_{2013-Q3}$0.02 0.67 -0.20 0.23 -0.09 $0.87^{**}$ 0.28 0.46 (se) (0.23) (0.40) (0.54) (0.39) (0.27) (0.25) (0.31) (0.21) $ATT_{2013-Q4}$ 0.22 $1.61^{**}$ 0.18 0.25 -0.44 $0.98^{**}$ 0.39 0.22 (se) (0.22) (0.50) (0.51) (0.37) (0.28) (0.28) (0.32) (0.19) $ATT_{2014-Q1}$ 0.06 0.94 0.83 0.25 0.39 $1.03^{**}$ $0.83^{*}$ 0.23 (se) (0.22) (0.43) (0.51) (0.38) (0.32) (0.29) (0.30) (0.20) $ATT_{2014-Q2}$ 0.24 $1.49^{**}$ 1.05 0.36 0.63 $1.00^{**}$ 0.77 0.36 (se) (0.21) (0.43) (0.53) (0.36) (0.28) (0.29) (0.34) (0.21) $ATT_{2014-Q3}$ $0.67^{**}$ $1.46^{**}$ 1.02 0.13 0.32 0.65 $0.95^{**}$ 0.44 (se) (0.21) (0.43) (0.52) (0.38) (0.30) (0.28) (0.30) (0.21) $ATT_{2014-Q4}$ 0.40 $1.76^{**}$ 0.86 0.19 0.16 $0.98^{**}$ 0.78 0.17 (se) (0.24) (0.47) (0.53) (0.37) (0.32) (0.29) (0.32) (0.20)
Appendix D Robustness checks
D.1 Results when excluding Southern-Danish households
Table 11: The causal effect of the Danish fat tax on the average quarterly consumption and expenditure per household during the tax period (October 2011 to December 2012). Southern-Danish households located in South/West Jutland and Funen are excluded.
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Quarterly total weight consumed in gr/ml Danish fat tax (ATT) -51.848 56.607 $-417.038^{**}$ 6.714 -116.145 118.680 -58.025 18.472 (se) (39.344) (87.142) (122.812) (74.980) (51.767) (98.194) (37.582) (40.566) % change -9.39% 3.72% -15.42% 0.50% -10.99% 6.93% -11.39% 3.30% obs. 23,844 34,792 26,029 33,376 25,232 32,761 25,186 30,678 nr. of Danish hh. 1,690 1,892 1,841 1,700 1,813 1,647 1,752 1,520 nr. of North.-German hh. 553 1,616 698 1,596 627 1,614 660 1,409 Panel B: Quarterly total expenditure in Euro cents Danish fat tax (ATT) -25.626 $227.535^{***}$ $-294.203^{*}$ $63.645^{*}$75.550 $79.356^{**}$42.121 22.480 (se) (28.542) (41.428) (110.695) (22.569) (31.786) (23.360) (37.458) (11.112) % change -7.13% 24.07% -12.52% 13.16% -13.10% 15.28% -7.78% 13.42% obs. 23,844 34,792 26,029 33,376 25,232 32,761 25,186 30,678 nr. of Danish hh. 1,690 1,892 1,841 1,700 1,813 1,647 1,752 1,520 nr. of North.-German hh. 553 1,616 698 1,596 627 1,614 660 1,409 Panel C: Quarterly total amount (in packages) consumed Danish fat tax (ATT) -0.058 0.278 -0.370 -0.272 -0.438 0.406 0.005 0.014 (se) (0.164) (0.334) (0.383) (0.262) (0.205) (0.215) (0.207) (0.148) % change -3.09% 4.64% -8.43% -8.41% -13.95% 11.19% 0.19% 1.13% obs. 23,844 34,792 26,029 33,376 25,232 32,761 25,186 30,678 nr. of Danish hh. 1690 1,892 1,841 1,700 1,813 1,647 1,752 1,520 nr. of North.-German hh. 553 1,616 698 1,596 627 1,614 660 1,409 Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors are clustered at the household level. % change compares the estimated tax average to the pre-tax sample mean for Danish households. The estimates are obtained using the doubly robust DID approach, controlling for age of head of household, number of children under 15, gender, ISCED, income level, and household size.
D.2 Results under unconditional parallel trends for the weight outcome
Figure 31: Bacon.
Figure 32: Butter.
Figure 33: Cheese.
Figure 34: Cream.
Figure 35: Liver sausage.
Figure 36: Margarine.
Figure 37: Salami.
Figure 31: Bacon.
Figure 32: Butter.
Figure 33: Cheese.
Figure 34: Cream.
Figure 35: Liver sausage.
Figure 36: Margarine.
Figure 37: Salami.
Figure 38: Sour cream.
Figure 39: Average quarterly weight consumed in gr/ml for the different product categories. The figures depict the trends when unconditional parallel trends is assumed and provide insights into the strength of the influence of the household characteristics on the differences in trends between Northern-German and Danish households. The trends are similar to those observed under conditional parallel trends (see Appendix C), indicating that the household characteristics are not strongly influencing differences in consumption trends between Danish and Northern-German households.
Appendix E Supplemental material for the heterogeneity analysis
(a) Average quarterly weight consumed (gr/ml).
(b) Average quarterly expenditure in Euro cents.
(c) Average quarterly number of packages.
Figure 40: Doubly robust DID estimates of the causal effect of the Danish fat tax during the post-tax period (January 2013 to December 2014) compared to the last taxed quarter (October 2012 to December 2012) for low, medium and high-income households. Low-income households are defined as households with very low income or low income. High-income households are defined as households with very high income or high income. 95% confidence intervals in red.